On January 1, 2017, the dental partnership of Angela, Diaz, and Krause was formed when the partners contributed $30,000, $58,000, and $60,000, respectively. Over the next three years, the business reported net income and (loss) as follows:

2017 $ 70,000
2018 42,000
2019 (25,000 )

During this period, each partner withdrew cash of $15,000 per year. Krause invested an additional $5,000 in cash on February 9, 2018.

At the time that the partnership was created, the three partners agreed to allocate all profits and losses according to a specified plan written as follows:

  • Each partner is entitled to interest computed at the rate of 10 percent per year based on the individual capital balances at the beginning of that year.
  • Because of prior work experience, Angela is entitled to an annual salary allowance of $12,000 per year and Diaz is entitled to an annual salary allowance of $9,000 per year.
  • Any remaining profit will be split as follows: Angela, 20 percent; Diaz, 40 percent; and Krause, 40 percent. If a net loss remains after the initial allocations to the partners, the balance will be allocated: Angela, 30 percent; Diaz, 50 percent; and Krause, 20 percent.

Determine the ending capital balance for each partner as of the end of each of these three years.
Answer #1
Date Particulars Angela Diaz Krause Total
January 1 2017 Opening Balance 30000 58000 60000 148000
December 31 2017 Interest 3000 5800 6000 14800
December 31 2017 Salary 12000 9000 0 21000
December 31 2017 Profit (70000-14800-21000 = 34200) 6840 13680 13680 34200
December 31 2017 Withdrawl -15000 -15000 -15000 -45000
December 31 2017 Balance 36840 71480 64680 173000
February 9 2018 Deposit 5000 5000
December 31 2018 Interest 3684 7148 6468 17300
December 31 2018 Salary 12000 9000 0 21000
December 31 2018 Profit (42000-17300-21000 = 3700) 740 1480 1480 3700
December 31 2018 Withdrawl -15000 -15000 -15000 -45000
December 31 2018 Balance 38264 74108 62628 175000
December 31 2019 Interest 3826.4 7410.8 6262.8 17500
December 31 2019 Salary 12000 9000 0 21000
December 31 2019 Profit (-25000 - 17500 -21000 = -63500) -12700 -25400 -25400 -63500
December 31 2019 Withdrawl -15000 -15000 -15000 -45000
December 31 2019 Balance 26390 50119 28491 105000
Answer #2

nope last year is different 

answered by: Yr